Sorry if this isn't the right topic; I don't quite know what kind of problem this is. If a does NOT equal b and 1/x + 1/a = 1/b, then x= ? if anyone can give the answer and steps for getting it, that would be super, many thanks!
Answer by richwmiller(17219)

The topic is fine.
You need to get x on one side and everything else on the other
1/x + 1/a = 1/b
1/b-1/a=1/x
x(1/b-1/a)=1
x=ab/(a-b)